OpenStudy (anonymous):

Simplify the following expressions. the product-to-sum formulas: (1 / 2)[sin(u + v) + sin(u – v)]

OpenStudy (anonymous):

sin(u+v) = sinu cosv + sinv cos u sin(u-v) = sinu cosv - sinv cosu sin(u + v) + sin(u – v) = sinu cosv + sinv cos u + sinu cosv - sinv cosu =2sinu cosv (1 / 2)[sin(u + v) + sin(u – v)] = (1/2)(2sinu cosv) = sinu cos v